Gibson Manufacturing Company reported the following data regarding a product it manufactures and sells. The sales price is $44. Variable costs Manufacturing Selling $11 per unit 7 per unit Fixed costs Manufacturing Selling and administrative $160,000 per yeair $180,600 per year Required a. Use the per-unit contribution margin approach to determine the break-even point in units and dollars. b. Use the per-unit contribution margin approach to determine the level of sales in units and dollars required to obtain a profit of $156,000. c. Suppose that variable selling costs could be eliminated by employing a salaried sales force. If the company could sell 20,800 units, how much could it pay in salaries for salespeople and still have a profit of $156,000? (Hint: Use the equation method.)
Answer #1

A)

Contribution margin = sales - variable cost

Contribution margin = $44 - ($11 + $7) = $26

Break even unit = fixed cost / contribution per unit

= ($160000 + $180000)/$26

= 13078 ~ 13100 units

Sales in dollar = 13100 × 44 = $576400

B) required sales = ( fixed cost + desired profit) ÷ contribution per unit

= ($340000 + $156000)/ $26

= 19078 ~ 19100 units

Sales in dollar = 19100 × 44 = $84400

C) fixed cost of salaries = $44800

No of unit Sold = 20800

Total Contribution Required = Total Fixed Cost + profit required

Total Contribution Required = $160000 + $180000 + $156000 + Fixed Cost of Salaries

No of unit Sold = (Total Contribution Required)/(Sale Price - Variable cost per unit)

20800 = ($496000 + Fixed Cost of Salaries)/(44-18)

Fixed Cost of Salaries= $540800 - $496000

= $44800

I have given complete and detailed solution to your problem, in case of any query you can comment and

PLEASE DO GIVE POSITIVE RATING